OpenStudy (anonymous):

Which of the following constants can be added to x^2 - x to form a perfect square trinomial? a. 1/4 b. 1/2 c. 1

OpenStudy (anonymous):

c?

OpenStudy (cwrw238):

to do these divide coefficient of x by 2 then square the result

OpenStudy (cwrw238):

coefficient of x = -1

OpenStudy (anonymous):

1/2?

OpenStudy (cwrw238):

no -1 / 2 = -1/2 (-1/2)^2 = 1/4

OpenStudy (cwrw238):

one question - in triangle post - did you mean right-angled triangle? - thats what i assumed